OpenStudy (anonymous):

given a polynomial f(x) of degree 2, where f(x+1) -f(x) = 6x-8 and f(1)=26 then f(2) equals....? How do I solve this?

OpenStudy (anonymous):

if the polynomial is of degree 2 then write ax^2 + bx + c then write a(x +1)^2 + b(x+1) + c - (ax^2 + bx + c) = 6x - 8 since this corresponds to f(x+1) - f(x) then do a(1)^2 + b(1) + c = 26 since this corresponds to f(1) hopefully this reduces to a manageable system of equations that you can find the general equation quite easily with

hartnn (hartnn):

or you can just put x=1 in f(x+1) -f(x) = 6x-8 since you know f(1), you can find f(2), just by putting x=1 on both sides.

OpenStudy (anonymous):

do what @hartnn suggested, will save you tons of work

OpenStudy (anonymous):

so I will sub x for one on both sides. So it will be f(1+1)-f(1)=6(1)-8

hartnn (hartnn):

yes! go on....

OpenStudy (anonymous):

so will it work out to 2-1 = 6-8 ????? i just get confused at this part because i don't know what to do

OpenStudy (anonymous):

no remember it's f(2) - f(1) = 6(1) - 8

hartnn (hartnn):

f(1+1) = f(2) so, you have f(2) - f(1) =6-8 and you know f(1) = 26 just plug in and find f(2)

OpenStudy (anonymous):

??? could you explain it a different way...if you don't mind

hartnn (hartnn):